OpenStudy (anonymous):

How do I prove this by using trig identities: (2sin^2x-1)/(sinxcosx)/=tanx-cotx

OpenStudy (ghazi):

\[2 \sin^2 x-1= 2\sin^2x-(\sin^2x+\cos^2x)= \sin^2x-\cos^2x \]

OpenStudy (anonymous):

how does that equal to sin^2x-cos^2x

OpenStudy (ghazi):

sin^2 x + cos^2x=1

OpenStudy (anonymous):

sin^2x-cos^2x/sin^2xcosx..now what do i do

OpenStudy (ghazi):

\[\frac{ \sin^2 }{ \sin x *\cos x }- \frac{ \cos^2 x }{ sinx * \cos x }= tanx - \cot x\] clear?

OpenStudy (ghazi):

@sabika13 hope that helps you

OpenStudy (anonymous):

how doss sin^2/sinxcosx=tanx?

OpenStudy (anonymous):

does

OpenStudy (ghazi):

 how about now?